### v1

# The Strong Birthday Problem Revisited

Abstract
We revisit the Strong Birthday Problem (SBP) introduced in DasGupta2005 . The problem is stated as follows: what is the minimum number of people we have to choose so that everyone has a shared birthday with probability at least 1/2? We derive recurrence relations to compute the probability, and further show a nice connection to the associated Stirling numbers of the second kind to derive additional recurrences. We implement the recurrences using dynamic programming as well as compute the values using the combinatorial formula, and provide numerical results.
Keywords Birthday problem $\cdot$ Strong birthday problem $\cdot$ Stirling number of the second kind $\cdot$ Associated Stirling number of the second kind $\cdot$ Recurrence $\cdot$ Dynamic programming
1 Introduction
What is the probability that in a randomly chosen group of $n$ people, at least two will have the same birthday? This problem is famously known as the birthday problem (see Feller1968introductionV1 , page 33). This problem was first published by Richard von Mises in 1939 vonMises1939 , but generally attributed to Harold Davenport who came up with this about 12 years earlier, but did not publish it at the time.
A simple set up for the birthday problem assumes that the birthday of any person follows a uniform distribution over the days of the year with simplifying assumptions including but not limited to no selection bias, i.e., preference for a specific birthday or month, no interference between birthdays between two persons, e.g., the hospital delays one child birth by a few hours to accommodate another emergency child birth, and no twins, etc. It can be shown that we only need 23 people to achieve a the probability more than 1/2 that at least two people share the same birthday. This is somewhat counterintuitive, given that we have 365 days in a year (disregarding the leap year), we need 366 people to attain the probability of 1, by the pigeonhole principle. A simple variant of this problem generalizes to matching birthdays of couples, that is we need husband-husband and wife-wife birthday matches. It is easy to show that (replace 365 by $365^{2}$ in the original solution to the birthday problem), we need 431 couples to achieve a probability of 1/2 or more DasGupta2005 . Yet another variant asks for the minimum number of people required so that at least two have a birthday within $d$ adjacent days with probability at least 1/2 Naus1968 . It can be shown that for $d=1,2,3,4,5$, the corresponding minimum number of people needed are $n=23,14,11,9,8$, respectively.
Several variants and extensions have been proposed to the birthday problem over the past eight decades Naus1968 ; mosteller1962 ; Mathis1991 ; Nunnikhoven1992 ; DasGupta2005 ; inoue2008 . While many of them can be entertaining puzzles, their solutions, computationally efficient exact and approximation algorithms designed to compute those have been an active area of research. This stems from the fact that several real-world problems in the domains of computer science including hashing, cryptography bellare2004 , communication networks and load balancing in distributed systems eijkhout2019 , and problems in other domains including biology Green2024 and medicine Krawitz2015 , forensic science Kaye2013Beyond , statistics, social science and data science gold2024 , can be modeled as one of the variants of the birthday problem.
In this article, we consider the Strong Birthday Problem (SBP) introduced in DasGupta2005 . The problem asks to compute the minimum number of people we have to choose so that everyone has a shared birthday with probability at least 1/2. We provide both combinatorial solution (as we could not find one in the earlier publications) in Subsection 2.1, and also derive a recurrence relation to compute the probability in Subsection 2.2. Further, we show that the associated Stirling numbers of the second kind (see Subsection 2.3) has a nice connection to this problem, which we exploited to derive additional recurrence relations (see Subsection 2.4). Finally, we implement the recurrences using dynamic programming, and also computed the values using the combinatorial formula. We provide numerical results based on our implementation in Section 3, and discuss other extensions in Section 4. Throughout the paper, we also note a number of related problems that are extensions to SBP which can be considered in their right, and our solution easily generalizes to those instances.
2 The Strong Birthday Problem (SBP)
The strong birthday problem is stated as follows. In a group of $n$ people, what is the probability that each person has a shared birthday with some other person in the group? In our set up, we assume that the number of days in a year is $m$, and each person’s birthday is equally likely to be incident on one of the days of the year. That is, the probability that $i$th person’s birthday is incident on $j$th day of the year is $1/m$, for $1\leq i\leq n$ and $1\leq j\leq m$. While this is a simplifying assumption for the birthday related problems, in may be a reasonable assumption in other contexts where the problem can be recast as an instance of a variant of one of the birthday problems.
The paper DasGupta2005 gave a formula to compute the probability for SBP However, we could not find a derivation of it. Therefore, as a warm up, we give a derivation for the combinatorial formula for SBP.
2.1 A Combinatorial Formula for SBP
Let the number of people be $n$ and the number of days in a year be $m$. Let’s compute the number of ways in which $k$ singleton birthdays, i.e., a day with exactly one person’s birthday incident on it, can be formed. There are four steps in counting we need to do here.
  1. 1.
    There are $m\choose k$ ways of choosing $k$ birthdays out of $m$ days.
  2. 2.
    There are $n\choose k$ ways of choosing $k$ people out of $n$ people.
  3. 3.
    The above chosen $k$ birthdays form a bijection with the chosen $k$ people. The number of such bijections is $k!$.
  4. 4.
    Furthermore, the remaining nkn-k people can be assigned the remaining mkm-k days in $(m-k)^{n-k}$ ways.
Therefore, the number of ways of forming $k$ singleton birthdays is
$M_{m,n,k}$ $=$ ${m\choose k}{n\choose k}k!(m-k)^{n-k}.$
However, since the assignment of the remaining nkn-k people to the remaining mkm-k days can have singleton birthdays, we end up over-counting the number of singleton birthdays, which we successively correct for by using the inclusion-exclusion principle rosen2019 to finally obtain the expression $N_{m,n,k}$, for the number of ways of forming $k$ singleton birthdays:
$N_{m,n,k}$ $=$ $\sum_{j=k}^{n}(-1)^{j-k}{j\choose k}M_{m,n,j}=\sum_{j=k}^{n}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (1)
The number of ways $n$ people can be assigned $m$ days as birthdays is $m^{n}$, which is the size of the sample space. Therefore, the probability of having $k$ singleton birthdays is
$P\!\left(m,n,k\right)$ $=$ $\frac{1}{m^{n}}N_{m,n,k}=\frac{1}{m^{n}}\sum_{j=k}^{n}(-1)^{j-k}{j\choose k}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (2)
The above formula gives us the probability of having $k$ singleton birthdays, which itself can be interpreted as yet another variant of the birthday problem. Note that he remaining mkm-k days can have 0, 2 or more than 2 (at most nkn-k) birthdays incident on them.
If we use $k=0$ in the above formula, it gives the probability that each person in the group of $n$ people shares her/his birthday with someone in the group:
$P\!\left(m,n,0\right)$ $=$ $\frac{1}{m^{n}}\sum_{j=0}^{n}(-1)^{j}{m\choose j}{n\choose j}j!(m-j)^{n-j}.$ (3)
2.2 A Recurrence for SBP
For the strong birthday problem, we can derive a recurrence relation as follows. Let $T(j,k,n,m)$ denote the number of ways in which $n$ birthdays are distributed over $m$ possible days in a year so that there are $j$ days with 2 or more birthdays incident on them and $k$ days with singleton birthdays. Therefore, mjkm-j-k days are birthday-free. Clearly, $n\geq 2j+k$. There are three mutually exclusive cases to consider in setting up a recurrence for $T(j,k,n,m)$, as given below.
  1. Case 1:
    (Add to days with 2 or more birthdays). $n$th person’s birthday is incident on one of the $j$ days with 2 or more birthdays. Then, there are T(j,k,n1,m)T(j,k,n-1,m) ways in which the first n1n-1 birthdays are distributed so that there are $j$ days with 2 or more birthdays incident on them and $k$ days with singleton birthdays. Therefore, the number of ways this can be done is jT(j,k,n1,m)jT(j,k,n-1,m).
  2. Case 2:
    (Add to singleton birthdays). $n$th person’s birthday is incident on one of the $k+1$ singleton birthdays. In this case, after $n$th person is included, the number of singletons birthdays goes from $k+1$ to $k$ and the number of days with 2 or more birthday goes from j1j-1 to $j$. Therefore, the number of ways this can be done is (k+1)T(j1,k+1,n1,m)(k+1)T(j-1,k+1,n-1,m).
  3. Case 3:
    (Create a Singleton birthday). $n$th person forms a singleton birthday. There are T(j,k1,n1,m)T(j,k-1,n-1,m) ways in which the first n1n-1 birthdays are distributed so that there $j$ days with 2 or more birthdays incident on them and k1k-1 days with singleton birthdays. One of the remaining mj(k1)=mjk+1m-j-(k-1)=m-j-k+1 days can be used to realize the $n$th person’s birthday. Therefore, the number of ways this can be done is (mjk+1)T(j,k1,n1,m)(m-j-k+1)T(j,k-1,n-1,m).
Hence the recurrence is given by
$T(j,k,n,m)$ $=$ $jT(j,k,n-1,m)$ (4)
$+(k+1)T(j-1,k+1,n-1,m)$
$+(m-j-k+1)T(j,k-1,n-1,m).$
The base cases for the recurrence come from the following simple observations. When $n<2j+k$, $T(j,k,n,m)=0$ by the set up of the problem. When $j=k=n=0$, $T(0,0,0,m)=1$, for all $m\geq 0$, as there is one way of doing such an assignment – by doing nothing. When $j<0$ or $k<0$, $T(j,k,n,m)=0$. Also, when $m<j+k$, $T(j,k,n,m)=0$.
Therefore, we can write the complete recurrence as
$T(j,k,n,m)$ $=$ $\begin{cases}0,&\text{if $j<0\lor k<0\lor n<2j+k$}\lor m<j+k;\\ 1,&\text{if $j=0\land k=0\land n=0$};\\ jT(j,k,n-1,m)\\ +(k+1)T(j-1,k+1,n-1,m)\\ +(m-j-k+1)T(j,k-1,n-1,m),&\text{otherwise.}\end{cases}$ (5)
The above recurrence can be computed using dynamic programming.
For the strong birthday problem, $k=0$, and $1\leq j\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor$, and the number of ways the birthdays are formed so that each day have either 2 or more birthdays or is birthday-free is then
$N$ $=$ $\sum_{j=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}T(j,0,n,m),$ (7)
and hence the probability of this happening is given by
$P\!\left(m,n\right)$ $=$ $\frac{1}{m^{n}}N=\frac{1}{m^{n}}\sum_{j=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}T(j,0,n,m).$ (9)
The above formula can be computed using dynamic programming. Since in the recurrence $j\in[0,n/2]$, $k\in[0,n]$ as it can grow during recursion due to the middle term in the recurrence, and $n$ ranges from 0 to $n$, the total number of states are $\operatorname{O}\left(n^{3}\right)$.Since the summation runs for $\operatorname{O}\left(n\right)$ time, the total time complexity is $\operatorname{O}\left(n^{4}\right)$ and space complexity is $\operatorname{O}\left(n^{3}\right)$. Here we assume that the complexity of addition and multiplication of large numbers is constant thereby we did not consider the bit complexity. Also, we assume that memoization and retrieval time is also constant.
2.3 Associated Stirling Numbers of the Second Kind
Stirling numbers of the second kind, denoted by $n\brace k$}, counts the number of ways to partition $n$ distinct objects $\left\{a_{1},\ldots,a_{n}\right\}$ into $k$ nonempty partitions (subsets). For example, ${4\brace 2}=7.$
We can write a recurrence for $n\brace k$} as follows (see graham94 page 259):
${n\brace k}$ $=$ $\begin{cases}0,&\text{if $(n=0\land k>0)\lor(n>0\land k=0)$};\\ 1,&\text{if $(n=0\land k=0)\lor(n>0\land k=1)$};\\ {n-1\brace k-1}+k{n-1\brace k},&\text{otherwise.}\end{cases}$ (10)
The $r$-associated Stirling number of the second kind, denote by ${n\brace k}_{\geq r}$ is the number of ways of partitioning $n$ distinct objects, $\left\{a_{1},\ldots,a_{n}\right\}$, into $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{r}\end{array}\!\!\right\rfloor$ unlabeled partitions where each partition has size at least $r\geq 2$. Note that this is a restriction over the Stirling number of the second kind where $r\geq 1$. For example, ${7\brace 2}_{\geq 3}=35$ and ${7\brace 2}_{\geq 2}=105$.
We can derive a recurrence for $n\brace k$} using the following insight. The $n$th object $a_{n}$ can be considered in two mutually exclusive ways as follows.
  1. Case 1:
    (Forms a partition). $n$th object’s insertion creates a partition of size $r$. The number of ways in which r1r-1 objects can be chosen from n1n-1 objects is $n-1\choose r-1$, to which the $n$th object is added. The remaining nrn-r objects can be partitioned into k1k-1 partitions of size at least $r$ in ${n-r\brace k-1}_{\geq r}$ ways. Therefore, the number of ways are ${n-1\choose r-1}{n-r\brace k-1}_{\geq r}$.
  2. Case 2:
    (Does not form a partition). $n$th objects does not form a new partition. Then, there are ${n-1\brace k}_{\geq r}$ ways of partitioning n1n-1 objects into $k$ partitions of size at least $r$. $n$th objects can go into any of these $k$ partitions. Therefore, the number of ways are $k{n-1\brace k}_{\geq r}$.
Therefore, the recurrence is given by
${n\brace k}_{\geq r}$ $=$ $k{n-1\brace k}_{\geq r}+{n-1\choose r-1}{n-r\brace k-1}_{\geq r}.$ (11)
The base cases for the recurrence come from the following simple observations: ${0\brace 0}_{\geq r}=1$, and ${n\brace k}_{\geq r}=0$ when $n\leq 0$ or $k\leq 0$ or $n<kr$. Also, ${n\brace n}_{\geq r}=1$.
Therefore, we can write the complete recurrence as:
${n\brace k}_{\geq r}$ $=$ $\begin{cases}0,&\text{if $(n\leq 0\land k>0)\lor(n>0\land k\leq 0)\lor(n<kr)$};\\ 1,&\text{if $n=0\land k=0$};\\ k{n-1\brace k}_{\geq r}+{n-1\choose r-1}{n-r\brace k-1}_{\geq r},&\text{otherwise.}\end{cases}$ (12)
For $r=1$, the above recurrence reduces to Stirling numbers of the second kind. For $r=2$, the above recurrence reduces to
${n\brace k}_{\geq 2}$ $=$ $k{n-1\brace k}_{\geq 2}+(n-1){n-2\brace k-1}_{\geq 2}.$
See Comtet1974 ; Howard1980 ; Zhao2008 ; Komatsu2015 for more on the associated Stirling numbers of the second kind and related identities. We will use the above-derived recurrence for the associated Stirling numbers of the second kind to formulate a recurrence for SBP.
2.4 A Recurrence for SBP using the Associated Stirling Numbers of the Second Kind
When the partitions are labeled, as in case of the birthday problem, the number of ways of putting $n$ labeled objects into $k$ labeled partitions such that each partition has at least $r$ objects in it can be obtained by considering all $k!$ permutations of the $k$ labels designating the partitions. Therefore,
$T_{r}(n,k)$ $=$ $k!{n\brace k}_{\geq r}$ (13)
$=$ $\begin{cases}0,&\text{if $(n\leq 0\land k>0)\lor(n>0\land k\leq 0)\lor(n<kr)$};\\ 1,&\text{if $n=0\land k=0$};\\ k!\left(k{n-1\brace k}_{\geq r}+{n-1\choose r-1}{n-r\brace k-1}_{\geq r}\right),&\text{otherwise.}\end{cases}$ (14)
This recurrence has a nice interpretation as one more variant of the birthday problem – how many ways are there so that $n$ people’s birthdays can be assigned to $m$ days of a year so that each day has $r\geq 2$ person’s birthdays assigned to it, and the corresponding probability problem can be stated accordingly.
For $r=2$, the above recurrence reduces to
$T_{2}(n,k)$ $=$ $k!{n\brace k}_{\geq 2}=k!\left(k{n-1\brace k}_{\geq 2}+(n-1){n-2\brace k-1}_{\geq 2}\right).$ (15)
Alternative Recurrence for SBP: For the strong birthday problem, we can derive an alternative recurrence using the above recurrences as follows.
The number of ways in which $n$ people can have birthdays out of $m$ possible days so that each birthday must be shared by at least 2 people has the following steps in counting. The number of ways of choosing $k$ birthdays out of $m$ days is $m\choose k$}, where $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor$. Each such choice of $k$ has $T_{2}(n,k)=k!{n\brace k}_{\geq 2}$ ways of distributing $n$ people into $k$ days so that each day has at least 2 person’s birthdays incident on it. Thus, the total number of ways is given by
$N$ $=$ $\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}{m\choose k}T_{2}(n,k)=\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}{m\choose k}k!{n\brace k}_{\geq 2}.$ (18)
Since the number of ways in which $n$ people can be assigned $m$ days as birthdays is $m^{n}$, which is the size of the sample space, the probability that each person shares her/his birthday with someone in the group of $n$ people is
$P\!\left(m,n,r=2\right)$ $=$ $\frac{1}{m^{n}}N=\frac{1}{m^{n}}\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}{m\choose k}k!{n\brace k}_{\geq 2}.$ (20)
We can have a simple generalization to the strong birthday problem when we ask the following question: What is the probability that in a group of $n$ people and $m$ days in a year, each day has at least $r\geq 2$ birthdays incident or is birthday-free? In other words, what is the probability that each person’s birthday is shared by at least r1r-1 other people for $r\geq 2$? The solution can be written straightforwardly, by observing that $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{r}\end{array}\!\!\right\rfloor$, and is given by
$P\!\left(m,n,r\right)$ $=$ $\frac{1}{m^{n}}\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{r}\end{array}\!\!\right\rfloor}{m\choose k}k!{n\brace k}_{\geq r}.$ (22)
The above formula can be computed using dynamic programming. The outer summation takes $\operatorname{O}\left(n/r\right)$ time, and there are $\operatorname{O}\left(n^{2}/r\right)$ number of unique subproblems. Computation of the binomial coefficient $m\choose k$} and $k!$ takes $\operatorname{O}\left(n/r\right)$ time with memoization. Again assuming that per state we use only constant time (by disregarding the bit complexity of large number addition and multiplication). Fixing $r=2$, the overall time complexity is $\operatorname{O}\left(n^{3}\right)$ and the space complexity can easily be shown to be $\operatorname{O}\left(n^{2}\right)$.
The probability computation using the recurrence in Equation 22 is has better time and space complexities than the recurrence in Equation 9 we derived in Subsection 2.2.
3 Implementation and Numerical Results
We have implemented the combinatorial formula (Equation 3) and the above recurrences in Equation 9 and Equation 22 to compute the minimum number of people $n$ required for different values of the number of days $m$ so that each person has a shared birthday with someone in the group with probability at least 1/2 (see row 2 of Table 1) and at least 0.999 (see row 3 of Table 1). The results matched in all three cases, thereby proving the validity of our approaches.
$m$ 10 50 100 200 364 365 366 400 500 1000
$n_{@prob\geq 0.5}$ 41 304 690 1541 3054 3064 3073 3399 4375 9528
$n_{@prob\geq 0.999}$ 112 665 1410 2975 5653 5669 5686 6253 7937 16619
Table 1: Minimum number of people $n$ required for a given $m$ so that everyone has a shared birthday with someone in the group with probability at least 1/2 (row 2) and at least 0.999 (row 3).
For $n>200$, since the recurrence in the Equation 9 took much longer than the recurrence in the Equation 22 due to its higher time and space complexities. So our results for $n>200$ are computed only using Equation 3 and Equation 22. We use an arbitrary precision arithmetic library so that large integer operations can be performed. Overall, while the recurrence in Equation 22 is efficient for practical computation, we believe that the recurrence that led to Equation 9 is more intuitive and can just be derived from a sound understanding of the problem statement.
4 Significance and Extensions
As discussed in DasGupta2005 , the strong birthday problem has applications in fingerprint matching in criminology and sociology. There can be more applications in areas like job scheduling, such as modeling the queues of customers in a retail shopping store, where we might want to estimate the number of customers so that each queue has size $r\geq 2$ with a certain probability threshold. Such estimation may help in opening more queues or merging queues.
In this article, we also discussed natural extensions of the strong birthday problem which are interesting in their own right. Fast approximation algorithm to compute the values of $n$ given $m$ and a probability threshold, that can scale to arbitrary large numbers is an interesting direction to pursue. Restrictions to the uniform probability assumption, and cases where each person has a different probability to be incident on a day of a year, may have interesting properties to explore.